OpenStudy (anonymous):

find the zero of the polynomial function y=x^2-3x+2

OpenStudy (anonymous):

x^2 - 3 x + 2 =0 x^2 - (2 * 3/2 x) + (3/2)^2 - (3/2)^2 + 2 =0 (x - 3/2)^2 - (3/2)^2 + 2 =0 (x - 3/2)^2 - 9/4 + 8/4 =0 (x - 3/2)^2 - 1/4 =0 (x - 3/2)^2 - (1/2)^2 =0 (x - 3/2 - 1/2) (x - 3/2 + 1/2) =0 (x - 2) (x - 1) =0 x = 2, x=1